# Open Hardware in Quantum Technology

### v1

@@ I Introduction
Figure 1: Overarching diagram of the Open Quantum Hardware steps and their interconnection. 1: Design phase, which can involve a loop between simulation of the Hamiltonian and electromagnetic (EM) simulations used to define the QPU design and layout. 2: Fabrication step, e.g., through a foundry. 3: Installation and bring-up, which includes testing and characterization; information collected at this step can be fed back to step 1 to modify designs. 4: Sustained operation, with overlaps (with step 3), and is composed of data acquisition (with control and calibration) and can involve interfacing with infrastructure to provide cloud access to the device or experiment.
@@ II Open quantum hardware today
Open quantum hardware encompasses open-source software that is used for designing, analyzing, building, controlling, and programming quantum chips, foundries which build quantum chips, and cloud-accessible labs and testbeds that provide alternative, non-commercially-driven access to them. In this section, we review the state of open quantum hardware along each of these categories. It should be noted that within each of these categories, there can exist multiple stacks – configurations of components which are all inter-operable with one another, but not necessarily with components used in a different stack. Ensuring inter-operability of components across stacks would help ensure a more robust and sophisticated ecosystem. In Table 1 we summarize the different categories, providing examples of existing open hardware projects. The particular choices of components is again meant to be representative.
@@
Processor Design DASQA [45, 46], KQCircuits [47], PainterQubits/Devices.jl, pyEPR [8], Qiskit Metal [48], QuCAT [49]
Simulation and diagnostics KQCircuits [47], Pulser [50], Qiskit Metal [48], QuTiP [34, 35], QuTiP-QIP [39], sc-qubits [38]
Control and data acquisition ARTIQ [7], Duke-ARTIQ [51], Qua111partially open-source [52], QCoDeS [53], QICK [54], Quantify [55], QubiC [56], qupulse [57], Sinara Open Hardware [58, 59]
Remotely Accessible Labs222excluding commercial providers Forschungszentrum Jülich through OpenSuperQ [60], Quantum Inspire [61]
Testing (testbeds) Lawrence Berkeley National Lab’s AQT, Sandia National Labs’ QSCOUT [62], Sherbrooke’s Distriq DevTeQ
@@
Table 1: OQH examples grouped by functionalities. The different functionalities in open quantum hardware and examples of representative projects: Processor design, §II.1; simulation and diagnostics, §II.2; and control and data acquisition, §II.2.4. We also list programs related to facilites in OQH: Remotely accessible labs (where here we exclude here the many commercial providers and cloud services, that are discussed in §II.4.1); testing (testbeds), §II.4.2; and fabrication (foundries), §II.4.3.
@@ II.1 Blueprints and software for hardware design
Finite Element Methods (FEM) simulations, as those shown in Figure 2, can be obtained with PyEPR using the energy participation ratio approach (EPR). EPR unifies the design of dissipation and Hamiltonians around a single concept — the energy participation, a number between zero and one — in a single-step electromagnetic simulation. After the FEM simulations have validated the qubit properties (and possibly its coupling to a cavity), one is ready to fabricate the SC QPU. To this end, a photomask is generated to place the SC qubits on the substrate of a processor. You can find a typical CAD drawing of a superconducting flip chip QPU and an example mask generated by KQCircuits [47] in Figure  3 and more information about KQCircuits in the frame below (Example 1).
Figure 2: Finite Element Methods (FEM) chip simulations Design and FEM simulations of a cavity (left) and qubit (right) modes, using PyEPR [8]. The contour plot of cavity modes displays the intensity of the electric field at 9.0 GHz. The Current-density magnitude of a transmon qubit, showing transmon pads connected by a Josephson junction [adapted from Ref. [8]].
@@
Example 1. HARDWARE DESIGN.
KQCircuits: KLayout Python library for integrated quantum circuit design.
KQCircuits [47] is an open-source Python library created by IQM, a full-stack quantum computing startup, for designing superconducting quantum circuits. It automates the layout and simulation part of the design process, outputting chip layouts and photomask files in OASIS or GDSII formats, standard formats for the specification of data structures of photomasks, and integrated circuit layout. The layout files are sent to a mask manufacturer to produce a physical mask which is then used for the fabrication process. As a part of the mask export process, KQCircuits also produces other files such as netlists, to help with design verification. While KQCircuits itself cannot be used to perform simulations, it can be used to export automated simulation scripts and files for popular electromagnetic field simulation toolkits, such as Ansys HFSS/Q3D [72], pyEPR [8] (Energy Participation Ratio simulation), Sonnet [73], and Elmer [74]. These simulations use the parameterized geometry as the physical mask layouts produced by KQCircuits. A comparison between KQCircuits and Qiskit Metal is provided in Table 2.
KQCircuits generates multi-layer 2-dimensional-geometry representing common structures in quantum processing units (QPU). An important part of KQCircuits are the definitions of parameterized geometrical objects, or “elements”, and a framework for easily defining new ones. Generating the designs using code makes it easy to quickly create different variations of designs and helps to avoid costly human errors. The combination of many elements into a full QPU design is made easier by features such as named reference points used for automatic positioning of elements relative to each other and Graphical User Interface (GUI) editing. Furthermore, KQCircuits includes a library of premade chips, many of which have been manufactured and used for testing at IQM. KQCircuits works on top of KLayout [11], an open-source layout design software that is mainly used for classical electronics. Advantageously, existing KLayout functionalities can be used for quantum hardware other than classical, while KQCircuits only needs to add features specific to QPU design. This connection with KLayout can also help to bring together the wider electronics open-source hardware community and open quantum hardware community.
@@ II.2.1 Data acquisition
The data acquisition sofware uses instrument drivers to form an abstraction layer to the several pieces of hardware that control a QPU. Parameters can also be used in higher levels of abstraction on top of instrument parameters to represent QPU device elements that are controlled and read via the hardware layer, such as a qubit readout circuit. The parameters are then used in automated measurement routines with the goal to characterize, calibrate or program a QPU or its different elements. This requires saving, analyzing and visualizing the data that is recorded by these instruments in a dataset. Most data acquisition software frameworks therefore also include data storage abstractions and data visualization tools.
In order to control the QPU elements using quantum programming instructions, the control signals need to be calibrated for optimal gate and read-out fidelity. As shown in Fig. 1, calibration values are obtained via device testing and calibration measurements using the data acquisition software, analysis routines and data storage in the Cloud. More details can be found in section II.2.4 on optimal control, calibration and characterization for high-fidelity quantum operations.
@@
Example 2. CONTROL AND DATA ACQUISITION
QUASAR: A Quantum Instruction Set Architecture
The QUASAR development started in 2017 and went through several iterations. It is a project defining a quantum Instruction Set Architecture (ISA) to interface software with hardware. The first version was a tightly-coupled extension that required significant changes to the processor micro-architecture. Later, the implementation moved towards a more decoupled approach. The current implementation is a RISC-V Rocket Core [101] extended with the QUASAR co-processor (note that RISC-V is a very popular open standard ISA in classical open hardware. RISC-V is based on the established reduced instruction set computer (RISC) principles, and the Rocket Core is a processor that implements RISC-V.) In contrast to classical instruction sets, quantum ISAs operate on different types of computations. Traditional general-purpose operations are supplemented with a set of basic quantum gates applied on direct-addressed qubits and/or qubit registers. The RISC-V core executes the main program; when a quantum instruction occurs at the fetch pipeline stage, it forwards it via the RoCC interface [102] to be decoded and executed by the QUASAR co-processor. Such an architectural solution allows the main core to continue performing computations while the quantum backend is generating the control pulses. Moreover, the RISC-V core can perform complex computations, such as floating-point arithmetic for phase estimation, and make conditional branching during algorithm execution based on the qubit measurement data. That makes the QUASAR implementation flexible enough to accommodate complex hybrid experiments for classical-quantum computations. The RISC-V Rocket core runs a Linux kernel that facilitates remote communication with the user host machine and provides additional functionalities for the quantum software stack.
@@ II.2.2 Pulse-level control with hardware integration
Example 3. CONTROL AND DATA ACQUISITION.
ARTIQ: Advanced Real-Time Infrastructure for Quantum physics.
The ARTIQ experiment control and data acquisition system was initiated in 2013 at the NIST Ion Storage Group, in partnership with M-Labs Ltd, to address the deficiencies observed with control systems developed in-house by physicists or based on existing commercial solutions. A key feature of the ARTIQ system is a high-level Python-based programming language that helps describe complex experiments, which is compiled and executed on dedicated FPGA hardware with nanosecond timing resolution and sub-microsecond latency. Using the abstractions provided by Python, ARTIQ has the capability to handle the entire control stack from quantum circuits to pulse-level programming. ARTIQ also supports connecting several FPGAs together and synchronizing their clocks in the sub-nanosecond regime, which greatly expands the input/output (I/O) scalability of the system.
Initial versions of the ARTIQ system ran on physicist-designed hardware based on FPGA development kits with custom I/O expansion boards. In order to improve the quality, availability and reproducibility of the hardware, the Sinara project [59] was started in collaboration with Warsaw University of Technology. The Sinara project developed a modular system with carrier FPGA cards (Kasli, Kasli-SoC) controlling various so-called Eurocard Extension Modules (EEMs) catering to the needs of each experiment – such as digital I/O, Analog to Digital Converter (ADC), Digital to Analog Converter (DAC), Direct Digital Synthesis (DDS), Phase-Locked Loop (PLL) synthesizer, AWG. There is ongoing work by Duke University on firmware [51] and by Warsaw University of Technology on hardware [58] to port ARTIQ to the RF-SoC platform, with similar capabilities to QICK. Combined with the ARTIQ software, the Sinara hardware has encountered substantial success, with almost a thousand quantum physics experiments relying on ARTIQ/Sinara systems. ARTIQ, including its gateware, the firmware, and the ARTIQ tools and libraries are licensed as LGPLv3+. The Sinara hardware designs are licensed under CERN OHL. This ensures that a user of ARTIQ or Sinara hardware designs obtains broad rights to use, redistribute, study, and modify them.
@@
To take advantage of the growing capabilities of recent RF-SoC platforms, QubiC was then ported to the XilinX ZCU216 platform [107], capable of direct generation of control pulses at higher frequencies (up to 10GHz, in the second Nyquist zone). Additionally, a novel Distributed Processor design was created and implemented to enable distributed decision-making and branching, such as mid-circuit measurement and feed-forward [108]. The different iterations of QubiC implementations are shown in Figure 5. To form a prototype full-stack quantum control system with increased potential for future experiments, QubiC and QUASAR (described above in II.2.1) have been integrated and applied to quantum computing experiments at AQT, LBNL [109].
Figure 5: QubiC, through the years: QubiC 1.0, implemented initially on XilinX VC707, in its chassis with auxiliary components (left), the Analog Front End chassis for QubiC 1.0 (middle), and QubiC 2.0 implemented on XilinX ZCU216 with a custom SMA fanout board. A customized Surface Mount Assembly (SMA) fan-out board has been designed and recently fabricated to fully utilize all the channels of the ZCU216 board. Separately, a low-cost DAC extension was developed for the VC707 platform, to meet the varying frequency needs of different superconducting QPU architectures. The QubiC team has recently demonstrated heterogeneous synchronization of two sets of VC707 boards with low-cost DACs and two ZCU216 platforms [107].
@@ II.2.3 Pulse-level simulation
Example 4. PULSE LEVEL SIMULATION. Pulser: Library for pulse-level/analog control of neutral atom devices.
The Pulser framework is an open-source software library for designing pulse sequences for neutral-atom QPUs [50]. In such devices, individual atoms are placed in arrays of microscopic traps [114, 115, 116] and their quantum state is manipulated through the application of laser or microwave pulses. Using Pulser, developers can control all the relevant physical parameters of the qubit register and the driving channels.
The central object in Pulser is called a sequence, which is linked to a device. The device contains information about the hardware constraints, such as the maximal number of qubits available, the minimal pairwise distance between qubits which can be reached or the number of lasers and the maximal amount of laser power available for each of them. These constraints are enforced upon the register, where the neutral-atom array is defined, and upon the pulses that are added to the sequence. Each pulse is defined by its phase, amplitude and detuning waveforms, and pulses are sequentially added to the channels that are available on the device. The resulting program can subsequently be sent to QPUs and executed on the hardware after a device-specific calibration-aware compilation step. In addition to providing an interface to hardware, Pulser includes a built-in emulator relying on QuTiP [34] that faithfully reproduces the hardware behavior. To this extent Pulser acts also as a simulator for QPU design.

### v2

# Open Hardware Solutions in Quantum Technology

@@ I Introduction
Figure 1: Overarching diagram of the Open Quantum Hardware steps and their interconnection. 1: Design phase, which can involve a loop between simulation of the Hamiltonian and electromagnetic (EM) simulations used to define the QPU design and layout. 2: Fabrication step, e.g., through a foundry. 3: Installation and bring-up, which includes testing and characterization; information collected at this step can be fed back to step 1 to modify designs. 4: Sustained operation, with overlaps (with step 3), and is composed of data acquisition (with control and calibration) and can involve interfacing with infrastructure to provide cloud access to the device or experiment. While these steps appear also in more general quantum hardware projects that are not open, their separation and standardization is a required feature to interconnect open frameworks.
@@ II Open quantum hardware today
Open quantum hardware encompasses open-source software that is used for designing, analyzing, building, controlling, and programming quantum chips, foundries which build quantum chips, and cloud-accessible labs and testbeds that provide alternative, non-commercially-driven access to them. In this section, we review the state of open quantum hardware along each of these categories. It should be noted that within each of these categories, there can exist multiple stacks – configurations of components which are all inter-operable with one another, but not necessarily with components used in a different stack. Ensuring inter-operability of components across stacks would help ensure a more robust and sophisticated ecosystem. In Table 1 we summarize the different categories, providing examples of existing open hardware projects and facilities. We group them by functionalities: different open-source and open-hardware projects are present in processor design, simulation and diagnostics, control and data acquisition. We also group facilities and organizations: remotely accesible labs, testbeds, and foundries.
In Table 2, we highlight how different projects and facilities (emphasized in italics in the table) are distributed on the main quantum technology architectures: atoms, ions, photonics, spins and SC qubits. It is visible that some architectures are populated with more projects, such as SC qubits and atoms. On the contrary, just a few tools exist for ions, with notable exceptions provided by ARTIQ and the Open Quantum Design foundation. Tools that have a strong cross-platform adoption are shown in the last row of Table 2. In the following sections we explore in greater depth how different tools have been developed for different functionalities and are used in various architectures.
@@
Projects Processor Design DASQA [45, 46], KQCircuits [47], PainterQubits/Devices.jl, pyEPR [8], Qiskit Metal [48], QuCAT [49]
Simulation and diagnostics KQCircuits [47], Pulser [50], Qiskit Metal [48], QuTiP [34, 35], QuTiP-QIP [39], sc-qubits [38], Strawberry Fields [51]
Control and data acquisition ARTIQ [7], Duke-ARTIQ [52], Qua111partially open-source [53], QCoDeS [54], QICK [55], Quantify [56], QubiC [57], Qudi [58] qupulse [59], Sinara Open Hardware [60, 61]
Facilities Remotely Accessible Labs222excluding commercial providers Forschungszentrum Jülich through OpenSuperQ [62], Quantum Inspire [63]
Testing (Testbeds) Lawrence Berkeley National Lab’s AQT, Open Quantum Design, Sandia National Labs’ QSCOUT [64], Sherbrooke’s Distriq DevTeQ, NQCC [65]
@@
Table 1: OQH examples grouped by functionality. The different functionalities in open quantum hardware and examples of representative projects: Processor design, §II.1; simulation and diagnostics, §II.2; and control and data acquisition, §II.2.4. We also list programs related to facilites in OQH: Remotely accessible labs (where here we exclude here the many commercial providers and cloud services, that are discussed in §II.4.1); testing (testbeds), §II.4.2; and fabrication (foundries), §II.4.3.
Atoms Ions Photonics Spins SC qubits
Projects Bloquade.jl, Labscript-qc, Pulser, QICK ARTIQ Perceval, Strawberry Fields QICK, Qudi, Qupulse ARTIQ, DASQA, KQCircuits, PyEPR, QICK, Qiskit Metal, QubiC, Sc-qubits
Facilities AQT, OpenSuperQ Open Quantum Design, QSCOUT LPS Collaboratory LPS Collaboratory, Quantum Inspire
Architecture-agnostic QCoDeS, Qibolab, Qua, Quantify, QuOCS, QuTiP-QIP, QCUP
Table 2: OQH examples grouped by architecture. The different open hardware projects, grouped by architectures (atoms, ions, photonics, spins, and SC qubits) The bottom row includes projects that are architecture-agnostics. Italics is used to highlight programs at facilities (e.g., instead of software projects).
@@ II.1 Blueprints and software for hardware design
Finite Element Methods (FEM) simulations, can be obtained with PyEPR using the energy participation ratio approach (EPR). EPR unifies the design of dissipation and Hamiltonians around a single concept — the energy participation, a number between zero and one — in a single-step electromagnetic simulation. After the FEM simulations have validated the qubit properties (and possibly its coupling to a cavity), one is ready to fabricate the SC QPU. To this end, a photomask is generated to place the SC qubits on the substrate of a processor. You can find a typical CAD drawing of a superconducting flip chip QPU and an example mask generated by KQCircuits [47] in Figure  2 and more information about KQCircuits in the frame below (Example 1).
@@
Example 1, Hardware Design: “KQCircuits: KLayout Python library for integrated quantum circuit design.” KQCircuits [47] is an open-source Python library created by IQM, a full-stack quantum computing startup, for designing superconducting quantum circuits. It automates the layout and simulation part of the design process, outputting chip layouts and photomask files in OASIS or GDSII formats, standard formats for the specification of data structures of photomasks, and integrated circuit layout. The layout files are sent to a mask manufacturer to produce a physical mask which is then used for the fabrication process. As a part of the mask export process, KQCircuits also produces other files such as netlists, to help with design verification. While KQCircuits itself cannot be used to perform simulations, it can be used to export automated simulation scripts and files for popular electromagnetic field simulation toolkits, such as Ansys HFSS/Q3D [75], pyEPR [8] (Energy Participation Ratio simulation), Sonnet [76], and Elmer [77]. These simulations use the parameterized geometry as the physical mask layouts produced by KQCircuits. A comparison between KQCircuits and Qiskit Metal is provided in Table 3. KQCircuits generates multi-layer 2-dimensional-geometry representing common structures in quantum processing units (QPU). An important part of KQCircuits are the definitions of parameterized geometrical objects, or “elements”, and a framework for easily defining new ones. Generating the designs using code makes it easy to quickly create different variations of designs and helps to avoid costly human errors. The combination of many elements into a full QPU design is made easier by features such as named reference points used for automatic positioning of elements relative to each other and Graphical User Interface (GUI) editing. Furthermore, KQCircuits includes a library of premade chips, many of which have been manufactured and used for testing at IQM. KQCircuits works on top of KLayout [11], an open-source layout design software that is mainly used for classical electronics. Advantageously, existing KLayout functionalities can be used for quantum hardware other than classical, while KQCircuits only needs to add features specific to QPU design. This connection with KLayout can also help to bring together the wider electronics open-source hardware community and open quantum hardware community.
@@ II.2.1 Data acquisition
The data acquisition sofware uses instrument drivers to form an abstraction layer to the several pieces of hardware that control a QPU. Parameters can also be used in higher levels of abstraction on top of instrument parameters to represent QPU device elements that are controlled and read via the hardware layer, such as a qubit readout circuit. The parameters are then used in automated measurement routines with the goal to characterize, calibrate or program a QPU or its different elements. This requires saving, analyzing and visualizing the data that is recorded by these instruments in a dataset. Most data acquisition software frameworks therefore also include data storage abstractions and data visualization tools. In order to control the QPU elements using quantum programming instructions, the control signals need to be calibrated for optimal gate and read-out fidelity. As shown in Fig. 1, calibration values are obtained via device testing and calibration measurements using the data acquisition software, analysis routines and data storage in the Cloud. More details can be found in section II.2.4 on optimal control, calibration and characterization for high-fidelity quantum operations.
@@
Example 2, Control and Data Acquisition: “QUASAR: A Quantum Instruction Set Architecture.” The QUASAR development started in 2017 and went through several iterations. It is a project defining a quantum Instruction Set Architecture (ISA) to interface software with hardware. The first version was a tightly-coupled extension that required significant changes to the processor micro-architecture. Later, the implementation moved towards a more decoupled approach. The current implementation is a RISC-V Rocket Core [104] extended with the QUASAR co-processor (note that RISC-V is a very popular open standard ISA in classical open hardware. RISC-V is based on the established reduced instruction set computer (RISC) principles, and the Rocket Core is a processor that implements RISC-V.) In contrast to classical instruction sets, quantum ISAs operate on different types of computations. Traditional general-purpose operations are supplemented with a set of basic quantum gates applied on direct-addressed qubits and/or qubit registers. The RISC-V core executes the main program; when a quantum instruction occurs at the fetch pipeline stage, it forwards it via the RoCC interface [105] to be decoded and executed by the QUASAR co-processor. Such an architectural solution allows the main core to continue performing computations while the quantum backend is generating the control pulses. Moreover, the RISC-V core can perform complex computations, such as floating-point arithmetic for phase estimation, and make conditional branching during algorithm execution based on the qubit measurement data. That makes the QUASAR implementation flexible enough to accommodate complex hybrid experiments for classical-quantum computations. The RISC-V Rocket core runs a Linux kernel that facilitates remote communication with the user host machine and provides additional functionalities for the quantum software stack.
@@ II.2.2 Pulse-level control with hardware integration
Example 3, Control and Data Acquisition: “ARTIQ: Advanced Real-Time Infrastructure for Quantum physics.” The ARTIQ experiment control and data acquisition system was initiated in 2013 at the NIST Ion Storage Group, in partnership with M-Labs Ltd, to address the deficiencies observed with control systems developed in-house by physicists or based on existing commercial solutions. A key feature of the ARTIQ system is a high-level Python-based programming language that helps describe complex experiments, which is compiled and executed on dedicated FPGA hardware with nanosecond timing resolution and sub-microsecond latency. Using the abstractions provided by Python, ARTIQ has the capability to handle the entire control stack from quantum circuits to pulse-level programming. ARTIQ also supports connecting several FPGAs together and synchronizing their clocks in the sub-nanosecond regime, which greatly expands the input/output (I/O) scalability of the system. Initial versions of the ARTIQ system ran on physicist-designed hardware based on FPGA development kits with custom I/O expansion boards. In order to improve the quality, availability and reproducibility of the hardware, the Sinara project [61] was started in collaboration with Warsaw University of Technology. The Sinara project developed a modular system with carrier FPGA cards (Kasli, Kasli-SoC) controlling various so-called Eurocard Extension Modules (EEMs) catering to the needs of each experiment – such as digital I/O, Analog to Digital Converter (ADC), Digital to Analog Converter (DAC), Direct Digital Synthesis (DDS), Phase-Locked Loop (PLL) synthesizer, AWG. There is ongoing work by Duke University on firmware [52] and by Warsaw University of Technology on hardware [60] to port ARTIQ to the RF-SoC platform, with similar capabilities to QICK. Combined with the ARTIQ software, the Sinara hardware has encountered substantial success, with almost a thousand quantum physics experiments relying on ARTIQ/Sinara systems. ARTIQ, including its gateware, the firmware, and the ARTIQ tools and libraries are licensed as LGPLv3+. The Sinara hardware designs are licensed under CERN OHL. This ensures that a user of ARTIQ or Sinara hardware designs obtains broad rights to use, redistribute, study, and modify them.
@@
To take advantage of the growing capabilities of recent RF-SoC platforms, QubiC was then ported to the XilinX ZCU216 platform [110], capable of direct generation of control pulses at higher frequencies (up to 10GHz, in the second Nyquist zone). Additionally, a novel Distributed Processor design was created and implemented to enable distributed decision-making and branching, such as mid-circuit measurement and feed-forward [111]. The different iterations of QubiC 1.0 and QubiC 2.0 [112] implementations are shown in Figure 4. To form a prototype full-stack quantum control system with increased potential for future experiments, QubiC and QUASAR (described above in II.2.1) have been integrated and applied to quantum computing experiments at AQT, LBNL [113].
Figure 4: QubiC, through the years: QubiC 1.0, implemented initially on XilinX VC707, in its chassis with auxiliary components (left), the Analog Front End chassis for QubiC 1.0 (middle), and QubiC 2.0 [112] implemented on XilinX ZCU216 with a custom SMA fanout board. A customized Surface Mount Assembly (SMA) fan-out board has been designed and recently fabricated to fully utilize all the channels of the ZCU216 board. Separately, a low-cost DAC extension was developed for the VC707 platform, to meet the varying frequency needs of different superconducting QPU architectures. The QubiC team has recently demonstrated heterogeneous synchronization of two sets of VC707 boards with low-cost DACs and two ZCU216 platforms [110].
@@ II.2.3 Pulse-level simulation
Example 4, Pulse Level Simulation: “Pulser: A library for pulse-level/analog control of neutral atom devices.” The Pulser framework is an open-source software library for designing pulse sequences for neutral-atom QPUs [50]. In such devices, individual atoms are placed in arrays of microscopic traps [118, 119, 120] and their quantum state is manipulated through the application of laser or microwave pulses. Using Pulser, developers can control all the relevant physical parameters of the qubit register and the driving channels. The central object in Pulser is called a sequence, which is linked to a device. The device contains information about the hardware constraints, such as the maximal number of qubits available, the minimal pairwise distance between qubits which can be reached or the number of lasers and the maximal amount of laser power available for each of them. These constraints are enforced upon the register, where the neutral-atom array is defined, and upon the pulses that are added to the sequence. Each pulse is defined by its phase, amplitude and detuning waveforms, and pulses are sequentially added to the channels that are available on the device. The resulting program can subsequently be sent to QPUs and executed on the hardware after a device-specific calibration-aware compilation step. In addition to providing an interface to hardware, Pulser includes a built-in emulator relying on QuTiP [34] that faithfully reproduces the hardware behavior. To this extent Pulser acts also as a simulator for QPU design.